Trace the curce
Y=(x-2)(x+1)^2
1
Expert's answer
2019-12-16T09:25:40-0500

1 The points, where the function value Y is equal zero: "x=2" and "x=-1" .

2 Consider intervals.

When "x > 2" the function value Y is positive.

When "-1< x < 2" the function value Y is negative.

When "x < -1" the function value Y is also negative.

3 Let's transform the function expression:

"Y = (x-2)(x+1)^2 = (x-2)(x^2 +2x +1) ="

"(x^3 + 2 x^2 +x) - 2x^2 -4x - 2= x^3 -3x - 2"

Now we find the first derivative of the function:

"Y' = ( (x-2)(x+1)^2 )' = (x^3 -3x - 2)' = 3 x^2 - 3"

So, the function is smooth and its derivative "Y'" is equal zero at poins "x=1" and "x=-1" .

Also, "Y'<0" when "-1 < x < 1" (on this interval the function value is decreasing), and "Y'>0" when "x<-1" and "x>1" (in these intervals the function value is growing). So, it changes its sign from negative to positive at "x = 1" and from positive to negative at "x = -1". That is why the function has extrema at "x=-1" (maximum) and at "x =1" (minimum).